3(r + 3) + 2(r - 4)
Answer by lenny460(1073)

3(r + 3) + 2(r - 4)

3 * r = 3r
3 * 3 = 9
2 * r = 2r
2 * -4 = -8
Therefore:
3r + 9 + 2r - 8
3r + 2r = 5r
9 - 8 = 1

Answer:
5r + 1


Lennox Obuong
Algebra Tutor